OpenStudy (anonymous):

I still haven received help for this question and I really need it. How do i find the x and y intercept of the function for me to graph

OpenStudy (acxbox22):

x intercept:when y is 0 y intercept:when x is 0

OpenStudy (anonymous):

\[\frac{ 2 }{ x^{2} - 2x -3}\]

OpenStudy (anonymous):

@acxbox22 how do i find x and y ?

OpenStudy (acxbox22):

for the y intercept plug in 0 for x and solve for y for the x intercept plug in 0 for y and solve for x

OpenStudy (acxbox22):

\[y=\frac{ 2 }{ x^2-2x-3 }\] \[y =\frac{ 2 }{ (0)^2-2(0)-3 }\]

OpenStudy (acxbox22):

that is for the y intercept

OpenStudy (anonymous):

In got finding the Y but not finding x

OpenStudy (acxbox22):

@AccessDenied

OpenStudy (acxbox22):

^^^

OpenStudy (acxbox22):

i am confused on x intercept as well

OpenStudy (accessdenied):

For the x-intercept, we set y equal to 0... 0 = 2 / (x^2 - 2x - 3) The key thing to note is that a fraction will be zero only where its \(numerator\) is zero. Essentially, we can just multiply both sides by the denominator and remove it: 0 = 2 <-- now we solve for x here... except there are no solutions! So there are no x-intercepts!

OpenStudy (acxbox22):

thanks i was confused because i kept on thinking that x=2

OpenStudy (anonymous):

X is equal to 0 not 2 for clarification

OpenStudy (acxbox22):

there is no xintercept